A 250g , 23-cm-diameter plastic disk is spun on an axle through its center by an electric motor. What torque must the motor supply to take the disk from 0 to 1700 rpm in 4.1s ?
Answer #1

m = 0.250 kg
R = d/2 = 0.115 m
w = 1700*2?/60 = 178.024 rad/sec
I =